OpenStudy (anonymous):

Find sin θ if cot θ = -2 and cos θ < 0

OpenStudy (anonymous):

??

OpenStudy (anonymous):

cos(t) / sin(t) = -2 cos(t) = -2sin(t) using sin(t) = sqrt(1-cos^2(t)) we get cos(t) = -2sqrt(1-cos^2(t)) square both sides : cos^2(t) = 4(1-cos^2(t)) 5cos^2(t) = 4 cos(t) = +\- sqrt(4/5) you know that cos(t) < 0 so cos(t) = - sqrt(4/5)

OpenStudy (anonymous):

okay. so